Please help me. I am having problem with these questions.
1. Find three consecutive integeres such that the sum of their squares is 77.
2. 16x^2 - 2x - 3 Factor completely
3. 2x^2 - x - 10 Factor each polynomial completely
thank you kindly for your help.

Answer by checkley71(8403)

X^2+(X+1)^2+(X+2)^2=77
X^2+X^2+2X+1+X^2+4X+4=77
3X^2+6X+5-77=0
3X^2+6X-72=0
3(X^2+2X-24)=0
3(X+6)(X-4)=0
X=6=0
X=-6 ANSWER.
X+4=0
X=4 ANSWER.
PROOF
4^2+5^2+6^2=77
16+25+36=77
77=77